# Knowledge Prompting: How Knowledge Engineers Use Large Language Models

### v1

@@ 6.3. Acknowledgements
Appendix A Appendix: A
@@
With this research, we aim to investigate how knowledge engineers and practitioners use LLMs to develop and maintain KGs, ontologies, and knowledge bases. We will not collect any personal data. The responses to the questions are fully anonymous. The collected data will be shared with Otter AI, an automated transcription service hosted on an AWS server in the US. Your data will be processed under the terms of UK data protection law (including the UK General Data Protection Regulation (UK GDPR) and the Data Protection Act 2018). Transcripts will be kept until the publication of the research.
@@ Introduction
  1. (1)
    What is your experience with KE and building KGs/ontologies?
  2. (2)
@@
Appendix B Appendix: B
Table 6. Codebook created by applying inductive and deductive thematic analysis.
@@ Appendix B Appendix: B
LLMs cannot perform all KE tasks and still need humans in the loop
@@
Appendix C Appendix: C
Table 7. KE and LLM tools and methodologies with description and references.
Tools/Methodologies Description Reference
KE
eXtremeDesign methodology A framework for pattern based ontology design (Blomqvist et al., 2016)
@@ Appendix C Appendix: C
NeOn methodology A scenario-based methodology that supports the collaborative and dynamic aspects of ontology development, including in distributed environments (Suárez-Figueroa et al., 2011)
@@
HermiT Reasoner A reasoner for ontologies that, given an OWL file, can determine if the ontology consistent, identify subsumption relationships between classes, and more (Glimm et al., 2014)
@@
OOPS A web-based tool, independent of the ontology development environment, used for detecting modelling errors (Poveda-Villalón et al., 2012)
@@
OAEI Aims at comparing ontology matching systems on precisely defined test cases which can be based on ontologies of different levels of complexity and use different evaluation modalities (Euzenat et al., 2011)
@@
OntoClean A methodology for validating the ontological adequacy of taxonomic relationships based on formal, domain-independent properties of classes (Guarino and Welty, 2002)
LLMs
ChatGPT A language model developed by OpenAI (OpenAI, 2024)
PaLM A transformer-based model developed by Google (Anil et al., 2023)
@@
Llama A family of transformer-based autoregressive causal language models developed by Meta (Touvron et al., 2023)
@@
mPLUG-Owl A multimodal model that integrates images, text, and other information for comprehension and to generate responses (Ye et al., 2023)
@@
InstructBLIP A vision-language model leveraging instruction tuning to achieve state-of-the-art performance (Dai et al., 2024)
@@
Claude A language model developed by Anthropic with use cases such as summarisation, Q&A, and coding (Anthropic, 2024)

### v3

@@ 6.3 Acknowledgments
Appendix A Supplementary Materials for Interview Study
@@
With this research, we aim to investigate how knowledge engineers and practitioners use generative AI to develop and maintain KGs, ontologies, and knowledge bases. We will not collect any personal data. The responses to the questions are fully anonymous. The collected data will be shared with Otter AI, an automated transcription service hosted on an AWS server in the US. Your data will be processed under the terms of UK data protection law (including the UK General Data Protection Regulation (UK GDPR) and the Data Protection Act 2018). Transcripts will be kept until the publication of the research.
@@ Introduction
  1. 1.
    What is your experience with KE and building KGs and ontologies?
  2. 2.
@@
Appendix B Codebook for Thematic Analysis
See Table 6.
Table 6: Codebook created by applying inductive and deductive thematic analysis.
@@ Appendix B Codebook for Thematic Analysis
LLMs cannot perform all KE tasks and still need human-in-the-loop
@@
Appendix C List of Tools and Methodologies
See Table 7.
Table 7: KE and generative AI tools and methodologies with description and references.
(a) Knowledge Engineering
Tools & Methodologies Description
eXtremeDesign methodology [17] A framework for pattern based ontology design
@@ Appendix C List of Tools and Methodologies
NeOn methodology [165] A scenario-based methodology that supports the collaborative and dynamic aspects of ontology development, including in distributed environments
@@
HermiT Reasoner [55] A reasoner for ontologies that, given an OWL file, can determine if the ontology consistent, identify subsumption relationships between classes, and more
@@
OOPS [137] A web-based tool, independent of the ontology development environment, used for detecting modelling errors
@@
OAEI [41] Aims at comparing ontology matching systems on precisely defined test cases which can be based on ontologies of different levels of complexity and use different evaluation modalities
@@
OntoClean [59] A methodology for validating the ontological adequacy of taxonomic relationships based on formal, domain-independent properties of classes
(b) Generative AI
Tools & Methodologies Description
ChatGPT [122] A language model developed by OpenAI
PaLM [10] A transformer-based model developed by Google
@@
Llama [168] A family of transformer-based autoregressive causal language models developed by Meta
@@
mPLUG-Owl [184] A multimodal model that integrates images, text, and other information for comprehension and to generate responses
@@
InstructBLIP [29] A vision-language model leveraging instruction tuning to achieve state-of-the-art performance
@@
Claude [11] A language model developed by Anthropic with use cases such as summarisation, Q&A, and coding
Appendix D Example of KG Cards
See Table 8.
Table 8: KG Cards for the CS-KG [34]. The CS-KG is accessed through https://scholkg.kmi.open.ac.uk/. The dumps include the updated versions of the KG. The methods used for the KG’s development, the data, and some statistics are presented in the CS-KG scientific publication [34]. The publication describes the first version of the KG.
(a) Knowledge
Card Description
Domain Computer Science
Multilinguality English
Timeliness The CS-KG includes statements and entities extracted from scientific papers in the period 2010-2021. The team plans to keep maintaining and updating CS-KG in the following several years. The team created a fully automatic pipeline that we will run every six months to produce new versions of CS-KG that will include recent papers from OpenAlex [34].
Range 3.9M entities are classified as Methods, 1.3M as Tasks, 450M as Materials, 215K as Metrics, and 4M are associated with the OtherEntity.
Ontology Design Artifacts The CS-KG ontology builds on top of SKOS (https://www.w3.org/2004/02/skos/) and PROV-O (https://www.w3.org/TR/prov-o/). The design of the object properties started from a set of 39 high level predicates (e.g., uses, analyzes, includes) produced by the knowledge graph generation pipeline (see Section 4.2). The team then associate specific domain and range constraints to them, which are used to drive and correct the automatic extraction process [34].
(b) Provenance
Card Description
Provenance Artifacts PROV-O. Each statement in CS-KG includes: provo:wasDerivedFrom, which provides provenance information and lists the MAG IDs (now OpenAlex IDs) of the articles from which the statement was extracted; provo:wasGeneratedBy, which provides provenance and versioning information of the tools used to detect the statement [34].
Dataset & Data Cards Scientific articles selected by considering all papers from 2010 to 2019 with at least 1 citation (as of December 2021) and all the papers in 2020-2021 period from the set of articles from MAG [174] associated with the Field of Study “Computer Scienc”. Since MAG has been decommissioned in 2021, the following versions will adopt OpenAlex (https://openalex.org/), which offers a comparable publication coverage [34].
Generative AI Usage No information
Crowdsourcing No use of crowdsourcing.
(c) Graph Construction
Card Description
Data preprocessing No information
Methodology Based on the website (https://scholkg.kmi.open.ac.uk/), the KG was generated by applying an automatic pipeline that extracts entities and relationships using four tools: DyGIE++, Stanford CoreNLP, the CSO Classifier, and a new PoS Tagger. CS-KG publication [34] presents the full pipeline. This pipeline was evaluated on a manually crafted gold standard yielding competitive results. It then integrates and filters the resulting triples using a combination of deep learning and semantic technologies in order to produce a high quality KG. The CS-KG publication [34] states that 1200 statements were selected for evaluation. The set was manually annotated by 3 senior computer science researchers. The Fleiss’ kappa agreement [45] between the annotators was 0.435, indicating a moderate agreement. The majority vote schema was employed to generate the gold standard. In order to show the advantage of the hybrid method that builds on top of multiple tools, the authors compared the CS-KG pipeline against DyGIEpp, OpenIE, PoST, and against the union of their results (DyGIEpp + OpenIE + PoST). Table 2 in [34] reports the results of the evaluation in terms of precision, recall, and f-measure. The CS-KG pipeline outperforms all the other tools yielding an overall f-measure of 0.76.
Curation No information
Schema The schema is described in https://scholkg.kmi.open.ac.uk/cskg/documentation.php and https://scholkg.kmi.open.ac.uk/cskg/ontology.
Modalities Text
(d) Quality
Card Description
Accuracy The CS-KG publication [34] in Sections 4.2 and 4.3 describe in detail the automatic methods used to handle the syntactic and semantic errors.
Coverage 67M statements, 24M entities, 219 relations (https://scholkg.kmi.open.ac.uk/)
Coherency The CS-KG publication [34] in Section 4.4 describes in detail the automatic methods used to handle the error triples produced. There is no mention for constrain violations or other errors.
Completeness No information
Succinctness (i) SPARQL endpoint support (https://scholkg.kmi.open.ac.uk/sparql/), (ii) human-readable labels are provided, (iii) URIs, and (iv) the contact information is available at https://scholkg.kmi.open.ac.uk.
(e) Licensing
Card Description
License CS-KG is licensed under a Creative Commons Attribution 4.0 International License.
Publisher Team: Danilo Dessì, Francesco Osborne, Diego Reforgiato Recupero, Davide Buscaldi, Enrico Motta. This work is a collaboration of Scholarly Knowledge Modeling, Knowledge Media Institute, The Open University, University of Cagliari, FIZ Karlsruhe, Université Sorbonne Paris Nord.
Funding No information
Policy & Regulation CS-KG is aligned with the initiative of the Knowledge Graph Construction W3C Community Group for producing benchmarks, resources, and tools to support the semi-automatic generation of knowledge graphs from documents.
(f) Transparency
Card Description
Data Balance Statistics for the first version of CS-KG are provided in Section 5 of the publication [34].
Explainability No information
Limitations The main limitation of CS-KG is that it was produced with a fully automatic methodology, so the specific statements are not revised by humans, as in manually crafted KGs. The authors also working on developing an entity linking tool for automatically mapping documents (e.g., articles, patents, educational material) to entities and statements in CS-KG. Finally, they plan to further extend the ontology and the entity typing process, in particular by providing a more granular categorization of entity types.
Fairness Metrics No information
Applications CS-KG can support a variety of intelligent services, such as advanced literature search, document classification, article recommendation, trend forecasting, hypothesis generation, and many others [34].